OpenStudy (anonymous):

Evaluate 7a + 3b if a = -4 and b = -2.

OpenStudy (mimi_x3):

7(-4) + 3(-2) = -34

OpenStudy (diyadiya):

7(-4)+3(-2) =-28-6 =-34

OpenStudy (anonymous):

(7*-4)*(3*-2)=-28+-6=-34

OpenStudy (anonymous):

yay go with jesusisrisen his answer is correct